OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: x + y = 6 x = -y + 2

OpenStudy (anonymous):

x+y=6 x=-y+2 substitute (-y+2) for this x value (-y+2) + y = 6 simplify 2 = 6

OpenStudy (anonymous):

not really equal

OpenStudy (anonymous):

The solution is (2, 4). The solution is (4, 2). There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

you said to use substitution

OpenStudy (anonymous):

if you notice the question is wrong x+y = 6 x+y=2

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

no sol other wise :)

OpenStudy (anonymous):

You can see that these lines are parallel: